### v2

# VLMs Play StarCraft II: A Benchmark and Multimodal Decision Method

Abstract
We introduce VLM-Attention, a multimodal StarCraft II environment that aligns artificial agent perception with the human gameplay experience. Traditional frameworks such as SMAC rely on abstract state representations that diverge significantly from human perception, limiting the ecological validity of agent behavior. Our environment addresses this limitation by incorporating RGB visual inputs and natural language observations that more closely simulate human cognitive processes during gameplay. The VLM-Attention framework consists of three integrated components: (1) a vision-language model enhanced with specialized self-attention mechanisms for strategic unit targeting and battlefield assessment, (2) a retrieval-augmented generation system that leverages domain-specific StarCraft II knowledge to inform tactical decisions, and (3) a dynamic role-based task distribution system that enables coordinated multi-agent behavior. Our experimental evaluation across 21 custom scenarios demonstrates that VLM-based agents powered by foundation models (specifically Qwen-VL and GPT-4o) can execute complex tactical maneuvers without explicit training, achieving comparable performance to traditional MARL methods that require substantial training iterations. This work establishes a foundation for developing human-aligned StarCraft II agents and advances the broader research agenda of multimodal game AI. Our implementation is available at https://github.com/camel-ai/VLM-Play-StarCraft2.
starcraft2, multi-agent, benchmark, environment
1 Introduction
StarCraft II has established itself as a cornerstone benchmark for artificial intelligence systems, particularly in the domain of multi-agent reinforcement learning (MARL). While environments such as SMAC and PySC2 have facilitated significant advances in AI capabilities, they primarily rely on abstract state representations that fundamentally differ from human perception and decision-making processes. This abstraction, while computationally efficient, creates a significant disconnect between how AI agents and humans interact with and understand the game environment.
Recent breakthroughs in Vision-Language Models (VLMs) have demonstrated remarkable capabilities in bridging the gap between computer vision and natural language understanding. These models excel at tasks requiring complex reasoning, strategic planning, and adaptation to novel scenarios without explicit training. However, their potential in real-time strategy games, particularly in environments requiring both tactical decision-making and multi-agent coordination, remains largely unexplored.
We present VLM-Attention, a novel StarCraft II environment that fundamentally reimagines how AI agents perceive and interact with the game world. Our framework makes three key innovations: First, we redesign the observation space to align with human cognitive processes, incorporating both RGB visual input and natural language descriptions. This multimodal approach enables agents to process battlefield information in a manner analogous to human players, facilitating more intuitive strategic decision-making. Second, we develop a comprehensive action space that supports complex military maneuvers, including unit targeting, formation control, and ability usage. This design allows for fine-grained tactical control while maintaining accessibility for high-level strategic planning. Third, we introduce an integrated agent architecture that combines the reasoning capabilities of VLMs with specialized mechanisms for multi-agent coordination. We incorporate self-attention mechanisms for strategic unit targeting, retrieval-augmented generation (RAG) for leveraging domain-specific knowledge, and dynamic role assignment for effective task distribution.
To validate our approach, we create a diverse set of 12 micromanagement scenarios that test various aspects of tactical decision-making and unit control. Our experimental results demonstrate that VLM-based agents, utilizing models such as Qwen-VL and GPT-4o, can effectively execute complex tactical maneuvers without requiring extensive training, a significant departure from traditional MARL approaches. The primary contributions of this work include:
  • A novel multimodal StarCraft II environment that aligns agent perception with human cognitive processes, featuring RGB visual input, natural language observations, and a comprehensive action space for complex unit control.
  • An integrated VLM-based agent architecture that combines self-attention mechanisms, retrieval-augmented generation, and dynamic role assignment for effective tactical decision-making.
  • Empirical evidence demonstrating the effectiveness of VLM-based agents in executing complex StarCraft II micromanagement tasks without explicit training, supported by extensive evaluation across 12 scenarios.
Our work establishes a foundation for developing more human-aligned StarCraft II agents and contributes to the broader goal of creating AI systems that can understand and interact with complex game environments in ways that more closely mirror human cognition. In the supplementary material, we provide the code implementation and video demos of our agent decision-making.
2 Related Work
Vision-Language Model Benchmarks
In recent years, various VLM benchmarks have emerged that target different domains. VLMbench pioneered compositional designs for robotic manipulation, introducing structured evaluation of vision-language reasoning in physical tasks (Zheng et al., 2022). MultiMedEval (Royer et al., 2024) further advanced this direction by establishing a comprehensive framework for medical VLMs across multiple modalities, emphasizing standardized evaluation protocols. These benchmarks demonstrate the potential of VLMs in handling complex, domain-specific tasks, inspiring our approach to designing a VLM-based framework for StarCraft II that emphasizes human-like perception and decision-making processes.
General-Purpose Agents
The development of general-purpose agents has made significant strides in computer interaction tasks. CRADLE (Tan et al., 2024) established a unified framework using screenshots as input and keyboard/mouse actions as output, demonstrating the feasibility of vision-based control. WebVoyager (He et al., 2024) extended this approach through a sophisticated multimodal understanding of web interfaces, while ShowUI (Lin et al., 2024) introduced specialized techniques for GUI interaction, including UI-guided visual token selection and interleaved vision-language-action streaming. AppAgent (Zhang et al., 2023; Li et al., 2024a) further refined these concepts through simplified action spaces for smartphone control. These advances in general-purpose agents provide valuable insights for our work, particularly in designing intuitive interfaces and action spaces for StarCraft II environments.
StarCraft II AI Research
The evolution of StarCraft II AI systems has seen multiple paradigm shifts. AlphaStar (Vinyals et al., 2019) achieved superhuman performance through a combination of imitation learning and multi-agent reinforcement learning, followed by numerous improvements (Mathieu et al., 2021; star Contributors, 2021; Han et al., 2020; Huang et al., 2023). SMAC (Samvelyan et al., 2019; Ellis et al., 2023) provided a standardized framework for research but relied on abstract state representations that diverged from human perception. Recent works have begun exploring language models in StarCraft II, with LLM Play SC2 (Ma et al., 2024), LLM-PySC2 (Li et al., 2024b), and LLM-SMAC (Deng et al., 2024a) demonstrating the potential of language models in game understanding and decision-making. SMAC-Hard (Deng et al., 2024b) further extended these challenges by introducing more complex scenarios. Our work builds upon these foundations while introducing a novel approach that combines VLMs with attention mechanisms, creating a more human-aligned interaction paradigm that bridges the gap between abstract state representations and natural human gameplay.
3 Environment Design
Figure 1: Observation Space of environment.
Table 1: Comparison of StarCraft II environments. While SMAC simplifies the interface for MARL training and PySC2 provides comprehensive API access, VLM-Attention adapts PySC2 for multimodal decision-making through RGB observations and structured unit tracking.
SMAC PySC2 VLM-Attention
Base Environment PySC2 SC2 API PySC2
Visual Rep. None Feature layers (32x32x13) RGB screenshots
Unit Info Basic stats (health, coords) Detailed features (50x7) Tracked units with IDs
Battle Modes PVE with little abilities Full game modes PvP, PvE with abilities
Design Focus MARL Full game control Multimodal decision
Advantages Simple, stable Comprehensive API Visual processing Unit tracking Intuitive control
We formalize our StarCraft II environment as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},P,R)$, where agents receive observations and take actions in discrete time steps.
3.1 State and Observation Space
The true environment state $s\in\mathcal{S}$ contains the complete game information. However, agents only receive partial observations $o\in\mathcal{O}$ consisting of:
$o_{t}=(I_{t},T_{t},U_{t}),$ (1)
where:
  • $I_{t}\in\mathbb{R}^{H\times W\times 3}$ is an RGB image observation of the current game state with configurable resolution $(H,W)$.
  • $T_{t}$ is a natural language description of the game state, including faction-specific unit status.
  • $U_{t}=\{u_{1},...,u_{n}\}$ represents the set of visible unit information, where each $u_{i}$ contains:
    $u_{i}=(id_{i},type_{i},pos_{i},attr_{i},status_{i}),$ (2)
    with $id_{i}$ being the unit identifier, $type_{i}$ the unit class, $pos_{i}\in\mathbb{R}^{2}$ the position, $attr_{i}$ the unit attributes, and $status_{i}$ the current status values.
3.2 Action Space
The action space $\mathcal{A}$ is defined as the union of three action types:
$\mathcal{A}=\mathcal{A}_{attack}\cup\mathcal{A}_{move}\cup\mathcal{A}_{ability},$ (3)
where:
  • Attack actions $a_{attack}\in\mathcal{A}_{attack}$ are defined as ordered pairs $(i,j)$ representing unit $i$ attacking unit $j$
  • Move actions $a_{move}\in\mathcal{A}_{move}$ can be either:
    • Grid-based: $(i,x,y)$ where $i$ is the unit ID and $(x,y)\in\{1,...,10\}^{2}$
    • Directional: $(i,d)$ where $d\in\{\text{UP},\text{RIGHT},\allowbreak\text{DOWN},\text{LEFT}\}$
  • Ability actions $a_{ability}\in\mathcal{A}_{ability}$ are triples $(I,\allowbreak ability\_type,target)$ where $target$ can be a position, unit ID, or null depending on the ability type
3.3 Reward Structure and Episode Termination
We employ a sparse reward structure defined as:
$R(s_{t})=\begin{cases}1&\text{if victory}\\ -1&\text{if defeat}\\ 0&\text{if draw or ongoing}\end{cases},$ (4)
where the episodes terminate under three conditions: Victory: All enemy units eliminated; Defeat: All allied units eliminated; Draw: Time limit (300 seconds) exceeded.
3.4 Battle Modes and Technical Features
Our environment supports both Player vs. Player (PvP) and Player vs. AI (PvE) modes, built on top of PySC2. We designed 12 specialized micromanagement scenarios covering essential tactical elements: multi-unit coordination, ability usage (e.g., Marine stutter-step, Stalker blink), advanced maneuvers (split, hit-and-run), and terrain utilization. The implementation features an automated unit annotation system and a grid system for precise spatial control, alongside observation wrappers that capture RGB screenshots and generate natural language descriptions.
Figure 2: VLM-Agent Architecture integrates VLM-Attention, a knowledge-enhanced decision system, and a dynamic role assignment framework.
4 VLM-Agent Architecture
Our VLM-agent architecture integrates three key components: a self-attention mechanism for strategic unit targeting, a knowledge-enhanced decision system, and a dynamic role assignment framework. This design enables effective battlefield analysis and coordinated unit control while maintaining alignment with human decision-making processes.
4.1 VLM-Attention Mechanism
We propose VLM-Attention, an integrated system that combines skill planning and tactical decision-making through vision-language models. The mechanism processes battlefield information through two key stages to emulate human-like attention and decision-making processes.
First, we introduce a VLM Planner that evaluates the battlefield situation and generates specific micro-management skill plans:
$S=\text{VLM}_{\text{plan}}(I,T,H)=\{s_{\text{primary}},s_{\text{secondary}}\},$ (5)
where the planner outputs a structured skill plan: {mdframed}[backgroundcolor=verylightgray, linecolor=black, linewidth=1pt] { “primary_skill”: { “name”: “Focus Fire”, “description”: “Concentrating damage on specific targets”, “steps”: [ “Select highest priority target”, “Command all units to attack the same target” ] }, “secondary_skills”: […] }
Based on the planner’s output, the system then performs strategic unit identification:
$A=\text{VLM}_{\text{detect}}(I)=\{a_{1},...,a_{n}\},$ (6)
where each annotation $a_{i}=(p_{i},c_{i},b_{i})$ consists of unit position $p_{i}$, unit class $c_{i}$, and bounding box $b_{i}$.
The strategic unit identification is then performed through a skill-aware natural language prompt:
$U_{\text{priority}}=\text{VLM}_{\text{analyze}}(I,T,H,A,Q,S),$ (7)
where:
  • $I$ is the current game screenshot
  • $T$ is the text state description
  • $H$ represents action history
  • $A$ is the set of unit annotations
  • $Q$ is the tactical analysis prompt generated based on the skill plan
  • $S$ is the current skill plan from VLM Planner
The VLM outputs its analysis in natural language, integrating both skill-based strategy and tactical considerations. For example, in a complex battlefield scenario with multiple Protoss units facing diverse Terran forces, the VLM analyzes the situation and identifies critical targets: {mdframed}[backgroundcolor=verylightgray, linecolor=black, linewidth=1pt, roundcorner=10pt] Unit: Marine_1 (Tag: 7)
Reason: Aligning with our Focus Fire strategy, this unit’s low health (45/45) makes it an ideal concentrated target. Swift elimination will reduce the enemy’s overall damage output.
Unit: Ghost_1 (Tag: 9)
Reason: Can severely impact our units with EMP or Snipe abilities, directly threatening our shield management strategy for the Archon. Early elimination is crucial for our planned skill execution.
4.2 Knowledge Integration through RAG
To enhance the VLM’s tactical understanding with domain expertise, we implement a RAG system that operates on the priority units identified by VLM-Attention. Given the priority unit set $U_{\text{priority}}\subseteq A$, we formulate the knowledge retrieval and integration process as:
$K(u)=\text{Retrieve}(c_{u})=\{s_{u},m_{u},t_{u}\}\quad\forall u\in U_{\text{priority}},$ (8)
where for each unit $u$ with class $c_{u}$, we retrieve a knowledge tuple $K(u)$ consisting of unit specifications $s_{u}$ (attributes, statistics), matchup data $m_{u}$ (counter relationships), and tactical insights $t_{u}$ (competitive usage patterns). The retrieved knowledge is then integrated with the current game state through a context-aware generation process:
$D=\text{VLM}_{\text{synthesize}}(I,T,H,U_{\text{priority}},\{K(u)\}),$ (9)
where $D$ represents the tactical decision guidance generated by combining the retrieved knowledge with the current game state representation.
For example, given a priority Marine unit $u\in U_{\text{priority}}$ identified by VLM-Attention, the system retrieves structured knowledge $K(u)$ including combat statistics (DPS: 9.8 (+1.6)), counter relationships (Strong against: [Hydralisk, Immortal, Marauder]), and matchup-specific insights. The system then generates tactical guidance such as:
{mdframed}
[backgroundcolor=verylightgray, linecolor=black, linewidth=1pt, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerleftmargin=10pt, innerrightmargin=10pt] “Marine unit analysis: Current configuration indicates optimal engagement against identified Hydralisk units. Recommend leveraging Stimpack timing window and maintaining Medivac support distance of 5 range units. Priority should be given to securing positional advantage given the unit’s base movement speed of 3.15.”
This knowledge-grounded synthesis enables the agent to make tactically sound decisions based on both immediate battlefield conditions and established strategic principles.
4.3 Dynamic Role Assignment and Task Distribution
In multi-agent environments, particularly in complex dynamic scenarios like StarCraft II, efficient task allocation is fundamental to system performance optimization. Drawing inspiration from human societies’ division of labor and cooperation, we propose a novel VLM-Agent architecture that implements a unified approach to role assignment and task distribution, capable of adapting to rapidly evolving battlefield conditions while maintaining strategic coherence. Formally, let $\mathcal{N}=\{1,...,N\}$ denote the set of agents and $\mathcal{Z}=\{z_{1},...,z_{m}\}$ represent the set of possible roles. The role assignment function $\phi:\mathcal{N}\rightarrow\mathcal{Z}$ maps each agent to a specific role, while considering the state space $\mathcal{S}$ and action space $\mathcal{,T}$ of the environment. This mapping is dynamically optimized through a utility function $U(\phi,s)$ that evaluates the effectiveness of the role given the current state $s\in\mathcal{S}$.
Figure 3: Dynamic Role Assignment through VLM-Agent architecture.
The role assignment process incorporates a sophisticated reflection mechanism $\mathcal{M}$ that continuously evaluates performance through three key components:
1) Performance Monitoring: A real-time evaluation system that tracks key performance indicators (KPIs), including resource utilization efficiency, combat effectiveness, and strategic objective completion rates.
2) Historical Analysis: A temporal memory buffer that maintains a record of previous role assignments and their outcomes, enabling the system to learn from past experiences and adapt its strategies accordingly.
3) Adaptive Optimization: An iterative refinement process that adjusts role assignments based on both immediate feedback and historical performance patterns.
Our framework leverages a VLM through a sophisticated multimodal fusion function:
$z_{i}=\text{VLM}_{\text{role}}(I,T,C),$ (10)
where the model processes visual inputs $I$, textual prompts $T$, and contextual information $C$ to generate contextual representations. The system employs an iterative prompting mechanism with self-reflection, allowing the model to:
  • Generate comprehensive tactical assessments incorporating battlefield state, unit compositions, and resource distributions.
  • Evaluate potential outcomes through chain-of-thought reasoning to identify optimal strategies.
  • Refine decisions through structured prompt templates that encode tactical principles and maintain strategic consistency.
Through this sophisticated prompt engineering framework, our VLM-Agent architecture achieves robust battlefield performance while maintaining interpretable decision-making processes. The system dynamically adapts to battlefield conditions by iteratively refining its prompts based on unit capabilities, environmental contexts, and emerging tactical situations. This approach effectively bridges the gap between abstract state representations and human-like strategic reasoning, while maintaining computational efficiency through structured prompt optimization.
The effectiveness of our approach is further enhanced by incorporating uncertainty estimation and risk assessment mechanisms, allowing the system to make more nuanced decisions in ambiguous situations. This is particularly crucial in scenarios where incomplete information or rapidly changing conditions require robust and adaptable decision-making capabilities.
4.4 Complete Pipeline
We formalize our decision-making pipeline as an iterative process that maps POMDP observations to actions through a series of VLM-based transformations. At each timestep $t$, given an observation $o_{t}=(I_{t},T_{t},U_{t})$ from the environment state $s_{t}\in\mathcal{S}$, our system generates actions through the Algorithm 1.
The system maintains a history buffer $H_{t}$ for states and actions, and processes each step at 2Hz, maximizing the trade-off between strategic depth and real-time responsiveness. The sparse reward structure $R(s_{t})$ guides the objective of achieving victory while eliminating enemy units.
Algorithm 1 VLM-Attention Decision Pipeline for StarCraft II
Input: StarCraft II environment $env$, History buffer size $H$
1:  Initialize environment and get initial observation $o_{0}=(I_{0},T_{0},U_{0})=env.reset()$
2:  Initialize history buffer $\mathcal{H}$, total reward $R=0$
3:  while $env$ is not terminated do
4:     // Stage 1: Micro-skill Planning
5:     Generate skill plan $S_{t}=\text{VLM}_{\text{plan}}(o_{t},\mathcal{H})$
6:     // Stage 2: Strategic Unit Analysis
7:     Detect units $A_{t}=\text{VLM}_{\text{detect}}(I_{t})$
8:     for each unit $u_{i}\in U_{t}$ do
9:        Parse unit info $(id_{i},type_{i},pos_{i},attr_{i},status_{i})$
10:     end for
11:     Identify priority units $U_{\text{priority}}=\text{VLM}_{\text{analyze}}(o_{t},S_{t})$
12:     // Stage 3: Knowledge Integration
13:     for each unit $u\in U_{\text{priority}}$ do
14:        Retrieve unit knowledge $K(u)=\text{Retrieve}(type_{u})$
15:     end for
16:     // Stage 4: Action Generation
17:     Initialize action set $a_{t}=\{\}$
18:     for each friendly unit $i$ do
19:        if $i$ should attack then
20:           Add $(i,j)\in\mathcal{A}_{attack}$ to $a_{t}$ for target unit $j$
21:        else if $i$ should move then
22:           Add $(i,x,y)\in\mathcal{A}_{move}$ or $(i,d)$ to $a_{t}$
23:        else if $i$ should use ability then
24:           Add $(i,\text{ability},\text{target})\in\mathcal{A}_{ability}$ to $a_{t}$
25:        end if
26:     end for
27:     // Execute action and update
28:     Get the reward and next observation: $r_{t},o_{t+1}=env.step(a_{t})$
29:     Update history buffer $\mathcal{H}$
30:     $R\leftarrow R+r_{t}$
31:     $o_{t}\leftarrow o_{t+1}$
32:     if Victory or Defeat or TimeLimit then
33:        break
34:     end if
35:  end while
36:  return total reward $R$
5 Experiment
Table 2: Performance evaluation of VLM-Attention across different StarCraft II scenarios, used Qwen-VL-Plus. Results show varying success rates across different unit compositions and tactical situations, with higher performance in simpler scenarios and challenges in more complex engagements.
Map Scenario Win Rate (%)
vlm_attention_1 75
pvz_task6_level3 25
3m_vlm_attention 75
2s3z_vlm_attention 25
2s_vs_1sc_vlm_attention 0
2m_vs_1z_vlm_attention 10
2c_vs_64zg_vlm_attention 0
2bc1prism_vs_8m_vlm_attention 0
3s_vs_3z_vlm_attention 10
8marine_1medvac_vs_2tank 0
6reaper_vs8zealot_vlm_attention 0
8marine_2tank_vs_zerglings_ banelings_vlm_attention 25
Average Win Rate 20.4
Table 3: Performance comparison of different VLM models on two representative StarCraft II scenarios. Results demonstrate the relative capabilities of various models in handling both complex multi-unit scenarios (vlm_attention_1) and simpler tactical engagements (3m_vlm_attention).
Model Win Rate (%)
vlm_attention_1 3m_vlm_attention
GPT-4-Turbo 87 79
GPT-4o 79 81
GPT-4o-mini 76 76
Qwen-VL-Plus 75 75
Table 4: Ablation study on vlm_attention_1 map by GPT-4-Turbo. Checkmarks indicate which components are enabled in each configuration. Results demonstrate the individual and combined effects of Role Assignment, VLM-Attention mechanism, and Retrieval-Augmented Generation (RAG) on system performance.
Role Assignment VLM-Attention RAG Win Rate (%)
87
71
65
70
24
50
26
20
5.1 Experimental Setup
We evaluate our VLM-Attention framework across 12 specialized StarCraft II micromanagement scenarios. These scenarios are carefully designed to test different aspects of tactical decision-making, ranging from basic unit control to complex multi-unit coordination with ability usage.
For the VLM models, we primarily use GPT-4-Turbo as our main model due to its superior performance, while also testing with GPT-4o, GPT-4o-mini, and Qwen-VL-Plus for comparative analysis. Our VLM agents are built on Camel [^1] https://github.com/camel-ai/camel (Li et al., 2023), which supports convenient multi-agent system design and training. The environment implementation is built on PySC2222https://github.com/google-deepmind/pysc2 (Vinyals et al., 2017), with the environment running at 2Hz to balance decision quality and real-time performance.
5.2 Component Analysis Results
Our ablation study investigates the contribution of each key component in our framework: VLM-Attention mechanism, RAG system, and Role Assignment module. The results demonstrate strong synergistic effects between these components, with the complete system achieving the highest performance (87% win rate on vlm_attention_1).
VLM-Attention proves to be the most crucial component, particularly in complex scenarios requiring sophisticated target selection and tactical positioning. When enabled alone, it improves win rates by 30-40% compared to the baseline, highlighting its effectiveness in strategic decision-making.
The RAG system provides essential domain knowledge support, resulting in more consistent and tactically sound decisions. Its integration improves win rates by 20-25%, especially in scenarios requiring specific matchup understanding. Role Assignment further enhances coordination between units, contributing a 15-20% performance improvement when combined with other components.
5.3 Model Performance Analysis
Our comparative analysis of different VLM models reveals important trade-offs between performance, speed, and cost. GPT-4-Turbo achieves the best performance with win rates up to 87% but faces significant latency and cost challenges in real-time applications. GPT-4o and GPT-4o-mini offer better efficiency while maintaining reasonable performance (76-81% win rates).
Qwen-VL-Plus, despite being the most cost-effective option, shows limitations in knowledge consistency and instruction following. However, its accessibility makes it suitable for large-scale experiments and prototyping.
5.4 Scenario-Specific Performance
Our system demonstrates varying effectiveness across different scenario types. In basic engagement scenarios like 3m_vlm_attention and simple multi-unit coordination (vlm_attention_1), the system achieves high win rates (75%). These scenarios primarily test fundamental tactical decision-making and basic unit control.
However, performance degrades significantly in scenarios requiring precise ability usage or complex movement control. Maps requiring sophisticated micro-management, such as 2s_vs_1sc_vlm_attention and 2c_vs_64zg_vlm_attention, result in notably lower win rates (0-10%). This performance gap reveals current limitations in handling high-precision tactical maneuvers and ability timing.
5.5 Limitations and Challenges
Through our experimental evaluation, we identified several key challenges:
Spatial Understanding
VLMs show limited comprehension of StarCraft II’s spatial dynamics, particularly in scenarios requiring sophisticated terrain utilization. This limitation is evident in maps like 2c_64zg_vlm_attention (Figure 8), where opportunities for high-ground advantage are often missed.
Real-time Control
The high-frequency decision-making requirements of micromanagement pose significant challenges. Models often struggle with temporal consistency, exhibiting issues like excessive target switching and suboptimal movement commands.
Knowledge Application
We observe inconsistencies in domain knowledge application, variable output formatting, and occasional unit misidentification. These issues are particularly pronounced in complex scenarios requiring precise ability usage and unit coordination.
These limitations suggest that while our framework shows promise in strategic decision-making, significant improvements are needed in handling scenarios requiring precise micro-management and real-time control.
5.6 Emergent Capabilities
Through our experiments, we observe several remarkable emergent capabilities that VLM-based agents demonstrate without explicit training:
Strategic Unit Prioritization
In complex scenarios, particularly in vlm_attention_1 (Figure 7), agents demonstrate sophisticated tactical understanding by consistently prioritizing high-threat units (e.g., focusing Ghosts before they can use critical abilities) and low-health targets to maximize damage efficiency. This prioritization emerges naturally from the VLM’s understanding of unit capabilities and battlefield dynamics.
Multi-Unit Synergy
The system exhibits emergent coordination patterns, particularly evident in the 8marine_2tank_vs_zerglings_banelings_vlm_attention scenario (Figure 5). Here, the agent autonomously develops a sophisticated combined-arms strategy:
  • Marines automatically form protective formations around Siege Tanks.
  • Tanks are positioned to maximize splash damage against Banelings while minimizing friendly fire.
  • Unit positioning dynamically adjusts based on enemy movement patterns.
Adaptive Tactical Responses: The agents demonstrate the ability to develop complex micro-management skills without explicit programming:
  • Focus Fire: In 3m_vlm_attention, agents consistently identify and eliminate low-health targets to reduce incoming damage.
  • Formation Control: As demonstrated in the 8marine_1medvac_vs_2tank (Figure 6) through intelligent unit splitting to minimize splash damage.
  • Area Control: In 2c_vs_64zg_vlm_attention (Figure 4), Colossi are automatically positioned to maximize their line damage potential.
  • Resource Management: Systematic optimization of unit health and energy across various scenarios.
Context-Aware Decision Making: The system demonstrates an understanding of broader tactical contexts beyond immediate unit interactions:
  • Automatically adjusts aggression levels based on unit composition advantages.
  • Considers terrain features in positioning decisions.
These emergent capabilities are particularly noteworthy as they arise from the VLM’s general understanding of tactical principles, rather than from explicit training on StarCraft II mechanics. This suggests that VLM-based approaches may offer a promising direction for developing more flexible and adaptable game AI systems that can reason about complex tactical situations in ways similar to human players.
6 Conclusion
VLM-Attention represents a significant advancement in creating more human-like StarCraft II agents. By aligning artificial agent perception and decision-making processes with human cognition, our framework provides a more natural and effective environment for developing and testing AI systems. Future work will focus on expanding the system’s capabilities and evaluating its effectiveness in more complex scenarios.
7 Impact Statement
This work advances the field of multimodal AI decision-making through the lens of real-time strategy games. While our primary contribution is methodological, we acknowledge several potential societal implications. The development of more human-aligned AI agents could enhance human-AI collaboration and improve AI system interpretability. However, advances in strategic decision-making capabilities also warrant careful consideration regarding dual-use applications. We believe our focus on human-centric design and transparent decision processes helps promote responsible AI development. Our framework primarily serves as a research tool for studying AI capabilities in controlled game environments, with minimal risk of direct negative societal impact.
Appendix A Map Details
Our environment features a diverse collection of 21 specialized maps, systematically categorized based on player count and ability usage capabilities. These maps originate from three primary sources: SMAC-based maps redesigned from the StarCraft Multi-Agent Challenge framework, original maps specifically designed for VLM-Attention evaluation, and selected scenarios adapted from the LLM-PySC2 framework [^3] https://github.com/NKAI-Decision-Team/LLM-PySC2.
Each map is meticulously designed to evaluate specific aspects of tactical proficiency and strategic decision-making:
  • Unit Control: Assessment of fundamental micromanagement capabilities
  • Multi-Unit Coordination: Evaluation of strategic control over heterogeneous unit compositions
  • Terrain Usage: Testing of positional awareness and environmental exploitation
  • Kiting: Assessment of dynamic hit-and-run tactical execution
  • Split: Evaluation of unit distribution strategies under enemy threats
  • Ability Usage: Testing of ability timing optimization and target prioritization
A.1 Single Player Maps (No Abilities)
Table 5: Single player maps without ability usage.
Map Name Unit Multi Terrain Kiting Split Mirror Units Source
Control Unit Usage Match
2c_vs_64zg_vlm_attention Player: 2 Colossi
Enemy: 64 Zerglings
SMAC
2m_vs_1z_vlm_attention Player: 2 Marines
Enemy: 1 Zealot
SMAC
2s_vs_1sc_vlm_attention Player: 2 Stalkers
Enemy: 1 Spinecrawler
SMAC
3s_vs_3z_vlm_attention Player: 3 Stalkers
Enemy: 3 Zealots
SMAC
6reaper_vs8zealot_vlm_attention Player: 6 Reapers
Enemy: 8 Zealots
NEW
8marine_1medvac_vs_2tank Player: 8 Marines, 1 Medivac
Enemy: 2 Siege Tanks
NEW
8marine_2tank_vs_zerglings _banelings_vlm_attention Player: 8 Marines, 2 Siege Tanks
Enemy: 35 Zerglings, 4 Banelings
NEW
2bc1prism_vs_8m_vlm_attention Player: 8 Marines
Enemy: 1 Warp Prism, 2 Photon Cannons
NEW
2s3z_vlm_attention Player: 2 Stalkers, 3 Zealots
Enemy: 2 Stalkers, 3 Zealots
SMAC
3m_vlm_attention Player: 3 Marines
Enemy: 3 Marines
SMAC
vlm_attention_1 Player: 1 Zealot, 1 Immortal, 1 Archon, 1 Stalker, 1 Phoenix
Enemy: 1 Marine, 1 Marauder, 1 Reaper, 1 Hellbat, 1 Medivac, 1 Viking (Assault), 1 Ghost, 1 Banshee
NEW
A.2 Single Player Maps (With Abilities)
Table 6: Single player maps with ability usage.
Map Name Unit Multi Terrain Kiting Split Ability Units Source
Control Unit Usage Usage
ability_map_8marine _3marauder_1medivac_1tank Player: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
Enemy: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
NEW
ability_8stalker_vs _8marine_3marauder _1medivac_tank_map Player: 8 Stalkers
Enemy: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
NEW
ability_map_8marine _3marauder _1medivac_1tank _vs_5stalker_2colossis Player: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
Enemy: 5 Stalkers, 2 Colossi
NEW
pvz_task6_level3 Player: 12 Stalkers, 1 Archon, 4 Sentries, 6 High Templars
Enemy: 64 Zerglings, 32 Banelings, 3 Ultralisks, 3 Queens
LLM-PYSC2
A.3 Two Player Maps (No Abilities)
Table 7: Two player maps without ability usage.
Map Name Unit Multi Terrain Kiting Split Mirror Units Source
Control Unit Usage Match
MMM_vlm_attention_two_players Player 1: 8 Marines, 3 Marauders, 1 Medivac
Player 2: 8 Marines, 3 Marauders, 1 Medivac
SMAC
vlm_attention_1_two_players Player 1: 1 Zealot, 1 Immortal, 1 Archon, 1 Stalker, 1 Phoenix
Player 2: 1 Marine, 1 Marauder, 1 Reaper, 1 Hellbat, 1 Medivac, 1 Viking (Assault), 1 Ghost, 1 Banshee
NEW
vlm_attention_2_terran_vs _terran_two_players Player 1: 1 Marine, 1 Marauder, 1 Reaper, 1 Hellbat, 1 Medivac, 1 Viking (Assault), 1 Ghost, 1 Banshee
Player 2: 1 Marine, 1 Marauder, 1 Reaper, 1 Hellbat, 1 Medivac, 1 Viking (Assault), 1 Ghost, 1 Banshee
NEW
A.4 Two Player Maps (With Abilities)
Table 8: Two player maps with ability usage.
Map Name Unit Multi Terrain Kiting Split Ability Units Source
Control Unit Usage Usage
ability_7stalker_vs _11marine_1medivac _1tank_map_2_players Player 1: 7 Stalkers
Player 2: 11 Marines, 1 Medivac, 1 Siege Tank
NEW
ability_8stalker_vs _8marine_3marauder _1medivac_tank_map_2_players Player 1: 8 Stalkers
Player 2: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
NEW
ability_map_8marine _3marauder_1medivac _1tank_2_players Player 1: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
Player 2: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
NEW
(a) Original RGB observation of battlefield situation
(b) Annotated unit positions with unit IDs and health status
(c) VLM-Attention’s strategic analysis highlighting prioritized targets and optimal attack vectors
Figure 4: Visualization of VLM-Attention’s decision-making process in a complex micro-management scenario. The system processes the raw visual input (a), analyzes unit positions and states (b), and makes tactical decisions (c). In this case, VLM-Attention identified Zergling_52 (Tag: 54) as a priority target due to its strategic position at [2,1], where attacking it would maximize area-of-effect damage to nearby clustered units. This decision demonstrates the system’s ability to not only identify low-health targets (5/35 HP) but also recognize opportunities for efficient damage distribution through Colossi’s line damage mechanic. Supporting this decision, the system also identified Zergling_1 (Tag: 3) and Zergling_2 (Tag: 4) as secondary priority targets due to their threatening positions at [1,1] and [0,1] ,respectively, enabling a comprehensive control strategy that combines focus fire with positional advantage.
(a) Initial state showing Marine/Tank positions
(b) VLM unit identification
(c) Priority targeting analysis
(d) Marine formation adjustment
(e) Coordinated focus fire execution
(f) Optimized Marine positioning
(g) Secondary target engagement
(h) Maintained spread formation
(i) Final engagement phase
Figure 5: Demonstration of VLM-Attention’s micro-management decision process in a Marine/Tank vs Baneling/Zergling scenario. The system implements a three-stage strategy: (1) Initial threat assessment prioritizing Banelings, (2) Coordinated unit positioning with Marines at safe distances ([0,1] and [0,2]) while maintaining focus fire capability, and (3) Sequential target elimination following the identified priority order of Banelings first, then supporting Zerglings. This showcases the system’s ability to combine Focus Fire, Formation Control, and Shield/Health Management skills in a complex tactical situation.
(a) Initial state analysis with two Siege Tanks identified as primary threats
(b) Execution of three-way split formation (3-3-2) with Medivac support
Figure 6: Visualization of VLM-Attention’s Marine split micro-management against Siege Tanks. The system first analyzes the battlefield state (a), identifying two Siege Tanks (Tag: 10, 11, Health: 160/160) as critical threats. Then it executes a sophisticated tactical response (b) by implementing a three-way Marine split formation (3-3-2) to minimize splash damage vulnerability while maintaining effective focus fire. The coordinated execution combines multiple micro-skills: Focus Fire for optimal damage distribution, Formation Control through strategic three-way positioning, and Health Management with Medivac support positioned at [0,2]. This demonstrates the system’s ability to balance multiple tactical objectives while maintaining combat effectiveness against superior firepower.
(a) Initial battlefield analysis with unit annotations
(b) Coordinated attack execution and positioning
Figure 7: VLM-Attention’s multi-type unit coordination in a complex Protoss vs Terran engagement. The system first performs threat assessment (a) by identifying critical targe,ts including a low-health Viking Assault (11/125 HP), energy-rich Ghost (56 energy), and support units like Medivac. Based on this analysis, it executes a coordinated attack plan (b) where each unit is assigned optimal targets: Zealot engages the weakened Viking, Phoenix provides air superiority against Medivac, Immortal focuses on armored targets, while the Archon maintains a strategic position at [7,7] for battlefield control. The decision-making process demonstrates VLM’s understanding of unit-specific attributes (health states, energy levels, armor types) and tactical synergies in a mixed-unit combat scenario. This showcases the system’s ability to handle complex micro-management tasks involving heterogeneous unit compositions without requiring explicit training for each unit combination.
Figure 8: Colossis stand in corner