### v5

# AVA: Attentive VLM Agent for Mastering StarCraft II

Abstract
We introduce Attentive VLM Agent (AVA), a multimodal StarCraft II agent that aligns artificial agent perception with the human gameplay experience. Traditional frameworks such as SMAC rely on abstract state representations that diverge significantly from human perception, limiting the ecological validity of agent behavior. Our agent addresses this limitation by incorporating RGB visual inputs and natural language observations that more closely simulate human cognitive processes during gameplay. The AVA architecture consists of three integrated components: (1) a vision-language model enhanced with specialized Multimodal Priority Inference mechanisms for strategic unit targeting and battlefield assessment, (2) a retrieval-augmented generation system that leverages domain-specific StarCraft II knowledge to inform tactical decisions, and (3) a dynamic role-based task distribution system that enables coordinated multi-agent behavior. The experimental evaluation in our proposed AVACraft environment, which contains 21 multimodal StarCraft II scenarios including both PvE and PvP settings, demonstrates that AVA powered by foundation models (specifically Qwen-VL and GPT-4o) can execute complex tactical maneuvers without explicit training, achieving comparable performance to traditional MARL methods that require substantial training iterations. Detailed case analyses further validate our approach, showing emergent tactical behaviors like focus-fire targeting and formation control that closely resemble professional human gameplay strategies. This work establishes a foundation for developing human-aligned StarCraft II agents and advances the broader research agenda of multimodal game AI. Our implementation is available at https://anonymous.4open.science/r/VLM-SMAC.
1 Introduction
StarCraft II has established itself as a cornerstone benchmark for artificial intelligence systems, particularly in the domain of multi-agent reinforcement learning (MARL). While environments such as SMAC and PySC2 have facilitated significant advances in AI capabilities, they primarily rely on abstract state representations that fundamentally differ from human perception and decision-making processes. This abstraction, while computationally efficient, creates a significant disconnect between how AI agents and humans interact with and understand the game environment. Recent breakthroughs in Vision-Language Models (VLMs) have demonstrated remarkable capabilities in bridging the gap between computer vision and natural language understanding. These models excel at tasks requiring complex reasoning, strategic planning, and adaptation to novel scenarios without explicit training. However, their potential in real-time strategy games, particularly in environments requiring both tactical decision-making and multi-agent coordination, remains largely unexplored.
We present Attentive VLM Agent (AVA), a novel StarCraft II multi-modal agent that fundamentally reimagines how AI agents perceive and interact with the game world. To evaluate the performance of AVA, we propose a multimodal StarCraft II environment, AVACraft, which makes three key innovations: First, we redesign the observation space to align with human cognitive processes, incorporating both RGB visual input and natural language descriptions. This multimodal approach enables agents to process battlefield information in a manner analogous to human players, facilitating more intuitive strategic decision-making. Second, we develop a comprehensive action space that supports complex military maneuvers, including unit targeting, formation control, and ability usage. This design allows for fine-grained tactical control while maintaining accessibility for high-level strategic planning. Third, we introduce an integrated agent architecture in AVA that combines the reasoning capabilities of VLMs with specialized mechanisms for multi-agent coordination. We incorporate Multimodal Priority Inference mechanisms for strategic unit targeting, retrieval-augmented generation (RAG) for leveraging domain-specific knowledge, and dynamic role assignment for task distribution.
We evaluate AVA in 12 diverse micromanagement scenarios that test various aspects of tactical decision-making and unit control. Our experimental results demonstrate that VLM-based agents, utilizing models such as Qwen-VL and GPT-4o, can effectively execute complex tactical maneuvers without requiring extensive training, a significant departure from traditional MARL approaches. The primary contributions of this work include:
  • We design a multimodal StarCraft II environment (AVACraft) that aligns agent perception with human cognitive processes, featuring RGB visual inputs, natural language observations, and a comprehensive action space for complex unit control.
  • We propose an integrated VLM-based agent architecture (AVA) that combines Multimodal Priority Inference mechanisms, retrieval-augmented generation, and dynamic role assignment for effective tactical decision-making.
  • The empirical evidence demonstrating the effectiveness of AVA in executing complex StarCraft II micromanagement tasks within AVACraft without explicit training, supported by extensive evaluation across 12 scenarios.
Our work establishes a foundation for developing more human-aligned StarCraft II agents and contributes to the broader goal of creating AI systems that can understand and interact with complex game environments in ways that more closely mirror human cognition. In the supplementary material, we provide the code implementation and video demos of our agent decision-making.
Figure 1: Observation Space of AVACraft environment.
2 Related Work
VLM Benchmarks and Agents
Recent advancements in Vision-Language Models (VLMs) have led to the development of various benchmarks and agent frameworks. VLMbench (Zheng et al., 2022) pioneered compositional evaluation for robotic manipulation tasks, while MultiMedEval (Royer et al., 2024) established standardized protocols for medical applications. For enhancing visual understanding, Set-of-Mark (Yang et al., 2023) introduced region-based prompting techniques that significantly improve GPT-4V’s grounding capabilities. Comprehensive evaluation frameworks like MME (Fu et al., 2024a) and Video-MME (Fu et al., 2024b) have further advanced the field by assessing both static and sequential visual reasoning across diverse scenarios. Building on these visual foundations, agent frameworks have emerged to enable practical applications. CRADLE (Tan et al., 2024) established a unified screenshot-to-action pipeline, while WebVoyager (He et al., 2024) and ShowUI (Lin et al., 2024) developed sophisticated techniques for web interface navigation and GUI interaction. AppAgent (Zhang et al., 2023; Li et al., 2024a) further refined these concepts through optimized action spaces for mobile environments.
StarCraft II AI Research
StarCraft II has served as a critical benchmark for artificial intelligence, particularly for multi-agent systems. AlphaStar (Vinyals et al., 2019) marked a milestone by achieving superhuman performance through a combination of imitation learning and MARL, inspiring numerous architectural improvements (Mathieu et al., 2021; star Contributors, 2021; Han et al., 2020; Huang et al., 2023). For standardized evaluation, the StarCraft Multi-Agent Challenge (SMAC) (Samvelyan et al., 2019; Ellis et al., 2023) provided a widely-adopted framework, though it primarily relied on abstract state representations that diverge from human perception. Recent advancements in this space include SMACv2, which introduced procedurally generated scenarios requiring adaptive closed-loop policies, and SMAC-Hard (Deng et al., 2024b), which extended the challenge with more complex tactical scenarios. The integration of language models with StarCraft II has been explored in works such as LLM Play SC2 (Ma et al., 2024), LLM-PySC2 (Li et al., 2024b), and LLM-SMAC (Deng et al., 2024a), demonstrating the potential of language-based reasoning for strategic game understanding.
VLMs for Complex Gaming Environments
Despite progress in both fields, StarCraft II presents unique challenges for VLMs due to its requirements for real-time decision-making and multi-agent coordination. Our work bridges this gap by introducing AVA, which leverages VLMs’ visual understanding while aligning agent perception with human gameplay experience, establishing a foundation for human-aligned StarCraft II agents without relying on abstract state representations.
3 AVACraft Environment Design
Previous StarCraft II AI research environments present significant limitations for human-aligned agent development. SMAC and SMACv2 employ abstract feature representations that create a perception gap between AI and human players, modifying unit attributes and using ”cheat mode” mechanics that deviate from actual gameplay. While SMACv2’s randomization tests generalization capabilities, these scenarios often fail to reflect competitive StarCraft II compositions or tactics employed in professional gameplay. The details of these limitations is provided in Appendix B.
Our AVACraft environment addresses these issues by processing RGB visual input and natural language observations, enabling pretrained Vision-Language Models (VLMs) to make tactical decisions with explicit reasoning that aligns with human perception and gameplay mechanics. And the map details can be found in D. We formalize our AVACraft environment as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},P,R)$, where agents receive observations and take actions in discrete time steps.
3.1 State and Observation Space
The true environment state $s\in\mathcal{S}$ contains the complete game information. However, agents only receive partial observations $o\in\mathcal{O}$ consisting of $o_{t}=(I_{t},T_{t},U_{t})$, where $I_{t}\in\mathbb{R}^{H\times W\times 3}$ is an RGB image observation of the current game state with configurable resolution $(H,W)$, capturing the visual battlefield representation. $T_{t}$ is a natural language description of the game state, including faction-specific unit status, battlefield conditions, and relevant tactical information. $U_{t}=\{u_{1},...,u_{n}\}$ represents the set of visible unit information, where each $u_{i}$ contains:
$u_{i}=(id_{i},type_{i},pos_{i},attr_{i},status_{i}),$ (1)
with $id_{i}$ being the unit identifier, $type_{i}$ the unit class, $pos_{i}\in\mathbb{R}^{2}$ the position coordinates, $attr_{i}$ the unit attributes (attack damage, armor, etc.), and $status_{i}$ the current status values (health, shields, energy, etc.).
Table 1: Comparison of StarCraft II environments. While SMAC and SMACv2 focus on reinforcement learning with abstract representations, AVACraft enables multimodal decision-making that aligns with human perception.
SMAC SMACv2 PySC2 AVACraft
Base Environment PySC2 PySC2 SC2 API PySC2
Visual Rep. None None Feature layers (32x32x13) RGB screenshots
Unit Info Basic stats (health, coords) Basic stats (health, coords) Detailed features (50x7) Tracked units with IDs
Battle Modes PVE with limited abilities Procedural scenarios Full game modes PvP, PvE with abilities
Design Focus MARL Generalization Full game control Multimodal decision
Advantages Simple, stable Randomization Generalization Comprehensive API Human-aligned perception Intuitive control
3.2 Action Space
The action space $\mathcal{A}$ is defined as the union of three distinct action types:
$\mathcal{A}=\mathcal{A}_{attack}\cup\mathcal{A}_{move}\cup\mathcal{A}_{ability},$ (2)
where attack actions $a_{attack}\in\mathcal{A}_{attack}$ are defined as ordered pairs $(i,j)$ representing unit $i$ attacking unit $j$, enabling precise targeting decisions. Move actions $a_{move}\in\mathcal{A}_{move}$ can be formulated in two complementary ways:
  • Grid-based: $(i,x,y)$ where $i$ is the unit ID and $(x,y)\in\{1,...,10\}^{2}$ represents a discrete position in the environment’s spatial grid.
  • Directional: $(i,d)$ where $d\in\{\text{UP},\text{RIGHT},\allowbreak\text{DOWN},\text{LEFT}\}$ facilitates relative movement commands.
Ability actions $a_{ability}\in\mathcal{A}_{ability}$ are triples $(i,\allowbreak ability\_type,target)$ where $target$ can be a position, unit ID, or null depending on the ability type, supporting unit-specific tactical capabilities.
3.3 Reward Structure and Episode Termination
We employ a sparse reward structure defined as:
$R(s_{t})=\begin{cases}1&\text{if victory}\\ -1&\text{if defeat}\\ 0&\text{if draw or ongoing}\end{cases},$ (3)
where the episodes terminate under three conditions: Victory: All enemy units eliminated; Defeat: All allied units eliminated; Draw: Time limit (300 seconds) exceeded. This approach provides clear performance signals while allowing flexible tactical execution.
3.4 Battle Modes and Technical Features
Our environment supports both Player vs. Player (PvP) and Player vs. AI (PvE) modes, built on top of PySC2. We designed 12 specialized micromanagement scenarios covering essential tactical elements: multi-unit coordination, ability usage (e.g., Marine stutter-step, Stalker blink), advanced maneuvers (split, hit-and-run), and terrain utilization. The implementation features an automated unit annotation system and a grid system for precise spatial control, alongside observation wrappers that capture RGB screenshots and generate natural language descriptions of battlefield conditions.
Figure 2: VLM-Agent Architecture integrates Multimodal Priority Inference mechanisms, a knowledge-enhanced decision system, and a dynamic role assignment framework.
4 Attentive VLM Agent Architecture
Our Attentive VLM Agent (AVA) architecture integrates three key components: a Multimodal Priority Inference mechanism for strategic unit targeting, a knowledge-enhanced decision system, and a dynamic role assignment framework. This architecture is designed to emulate human-like decision-making processes in complex StarCraft II scenarios.
4.1 Multimodal Priority Inference Mechanism
We propose a comprehensive system that combines structured skill planning and tactical decision-making through vision-language models. The mechanism processes battlefield information through two key stages to identify and prioritize strategic elements: First, we implement a VLM Planner that evaluates the battlefield situation and generates specific micro-management skill plans:
$S=\text{VLM}_{\text{plan}}(I,T,H)=\{s_{\text{primary}},s_{\text{secondary}}\},$ (4)
where the planner outputs a structured skill plan with primary and secondary tactical objectives: {mdframed}[backgroundcolor=verylightgray, linecolor=black, linewidth=1pt] { “primary_skill”: { “name”: “Focus Fire”, “description”: “Concentrating damage on specific targets”, “steps”: [ “Select highest priority target”, “Command all units to attack the same target” ] }, “secondary_skills”: […] }
Based on the planner’s output, the system performs precise unit identification and classification:
$A=\text{VLM}_{\text{detect}}(I)=\{a_{1},...,a_{n}\},$ (5)
where each annotation $a_{i}=(p_{i},c_{i},b_{i})$ consists of unit position $p_{i}$, unit class $c_{i}$, and bounding box $b_{i}$ for accurate spatial localization.
The critical Multimodal Priority Inference process then integrates visual features with tactical objectives through skill-aware natural language prompting:
$U_{\text{priority}}=\text{VLM}_{\text{analyze}}(I,T,H,A,Q,S),$ (6)
where $I$ is the current game screenshot capturing the visual battlefield state, $T$ is the text state description providing situational context, $H$ represents action history for temporal reasoning, $A$ is the set of unit annotations from the detection process, $Q$ is the tactical analysis prompt generated based on the skill plan, and $S$ is the current skill plan from the VLM Planner. The VLM outputs its analysis in structured natural language, integrating battlefield assessment with tactical prioritization. For example, in a complex scenario with multiple Protoss units facing diverse Terran forces, the system generates prioritized target assessments: {mdframed}[backgroundcolor=verylightgray, linecolor=black, linewidth=1pt, roundcorner=10pt] Unit: Marine_1 (Tag: 7)
Reason: Aligning with our Focus Fire strategy, this unit’s low health (45/45) makes it an ideal concentrated target. Swift elimination will reduce the enemy’s overall damage output.
Unit: Ghost_1 (Tag: 9)
Reason: Can severely impact our units with EMP or Snipe abilities, directly threatening our shield management strategy for the Archon. Early elimination is crucial for our planned skill execution.
4.2 Knowledge Integration through RAG
To enhance tactical decision-making with domain expertise, we implement a Retrieval-Augmented Generation (RAG) system that operates on the priority units identified through Multimodal Priority Inference. Given the priority unit set $U_{\text{priority}}\subseteq A$, we formulate the knowledge retrieval and integration process as:
$K(u)=\text{Retrieve}(c_{u})=\{s_{u},m_{u},t_{u}\}\quad\forall u\in U_{\text{priority}},$ (7)
where for each unit $u$ with class $c_{u}$, we retrieve a knowledge tuple $K(u)$ consisting of unit specifications $s_{u}$ (attributes, statistics), matchup data $m_{u}$ (counter relationships), and tactical insights $t_{u}$ (competitive usage patterns). The retrieved knowledge is then integrated with the current game state through a context-aware generation process:
$D=\text{VLM}_{\text{synthesize}}(I,T,H,U_{\text{priority}},\{K(u)\}),$ (8)
where $D$ represents the tactical decision guidance generated by combining the retrieved knowledge with the current game state representation. For example, given a priority Marine unit $u\in U_{\text{priority}}$ identified through our inference mechanism, the system retrieves structured knowledge $K(u)$ including combat statistics (DPS: 9.8 (+1.6)), counter relationships (Strong against: [Hydralisk, Immortal, Marauder]), and matchup-specific insights. The system then generates tactical guidance such as:
{mdframed}
[backgroundcolor=verylightgray, linecolor=black, linewidth=1pt, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerleftmargin=10pt, innerrightmargin=10pt] “Marine unit analysis: Current configuration indicates optimal engagement against identified Hydralisk units. Recommend leveraging Stimpack timing window and maintaining Medivac support distance of 5 range units. Priority should be given to securing positional advantage given the unit’s base movement speed of 3.15.”
4.3 Dynamic Role Assignment and Task Distribution
In multi-agent environments, efficient task allocation is fundamental to system performance optimization. Drawing inspiration from human societies’ division of labor and cooperation, we implement a unified approach to role assignment and task distribution that adapts to rapidly evolving battlefield conditions while maintaining strategic coherence. Formally, let $\mathcal{N}=\{1,...,N\}$ denote the set of agents and $\mathcal{Z}=\{z_{1},...,z_{m}\}$ represent the set of possible roles. The role assignment function $\phi:\mathcal{N}\rightarrow\mathcal{Z}$ maps each agent to a specific role, while considering the state space $\mathcal{S}$ and action space $\mathcal{A}$ of the environment. This mapping is dynamically optimized through a utility function $U(\phi,s)$ that evaluates the effectiveness of the role given the current state $s\in\mathcal{S}$.
Figure 3: Dynamic Role Assignment framework showing the evaluation and optimization process.
The role assignment process incorporates a sophisticated evaluation mechanism $\mathcal{M}$ that continuously assesses performance through three key components: (1) Performance Monitoring - a real-time evaluation system that tracks key performance indicators (KPIs), including resource utilization efficiency, combat effectiveness, and strategic objective completion rates; (2) Historical Analysis - a temporal memory buffer that maintains a record of previous role assignments and their outcomes, enabling the system to learn from past experiences and adapt its strategies accordingly; and (3) Adaptive Optimization - an iterative refinement process that adjusts role assignments based on both immediate feedback and historical performance patterns. Our framework leverages VLMs through a multimodal fusion function:$z_{i}=\text{VLM}_{\text{role}}(I,T,C)$, where the model processes visual inputs $I$, textual prompts $T$, and contextual information $C$ to generate contextual representations. The system employs an iterative analytical process with continuous evaluation. This process enables comprehensive tactical assessments incorporating battlefield state, unit compositions, and resource distributions; evaluation of potential outcomes through structured reasoning to identify optimal strategies; and refinement of decisions through parameterized prompt templates that encode tactical principles while maintaining strategic consistency.
Through this sophisticated prompt engineering framework, our architecture achieves robust battlefield performance while maintaining interpretable decision-making processes. The system dynamically adapts to battlefield conditions by iteratively refining its prompts based on unit capabilities, environmental contexts, and emerging tactical situations. This approach effectively bridges the gap between abstract state representations and human-like strategic reasoning, while maintaining computational efficiency through structured prompt optimization. The effectiveness of our approach is further enhanced by incorporating uncertainty estimation and risk assessment mechanisms, allowing the system to make more nuanced decisions in ambiguous situations. This is particularly crucial in scenarios where incomplete information or rapidly changing conditions require robust and adaptable decision-making capabilities.
4.4 Complete Pipeline
We formalize our decision-making pipeline as an iterative process that maps POMDP observations to actions through a series of VLM-based transformations. At each timestep $t$, given an observation $o_{t}=(I_{t},T_{t},U_{t})$ from the AVACraft environment state $s_{t}\in\mathcal{S}$, our system generates actions through the Algorithm 1 in Appendix C. The system maintains a history buffer $H_{t}$ for states and actions, and processes each step, maximizing the trade-off between strategic depth and real-time responsiveness. The sparse reward structure $R(s_{t})$ guides the victory while eliminating enemy units.
5 Experiments
We evaluate the Attentive VLM Agent (AVA) in the AVACraft environment to assess its tactical decision-making capabilities in StarCraft II micromanagement scenarios. Our experiments focus on two primary objectives: (1) quantifying AVA’s performance across diverse scenarios and vision-language models (VLMs), and (2) validating its alignment with human cognitive processes through expert evaluations. The implementation leverages the Camel framework (https://github.com/camel-ai/camel(Li et al., 2023) for multi-agent system design and PySC2 (https://github.com/google-deepmind/pysc2(Vinyals et al., 2017) for environment simulation, operating at 2Hz frequency to balance strategic decision depth and real-time responsiveness.
5.1 Quantitative Performance
We evaluated AVA across 12 micromanagement scenarios in AVACraft, designed to probe various tactical skills, including unit control, multi-unit coordination, and ability usage. These scenarios range from simple engagements (e.g., 3m) to complex battles requiring precise maneuvers (e.g., 2c_vs_64zg). We benchmarked four state-of-the-art VLM models—GPT-4-Turbo, GPT-4o, GPT-4o-mini, and Qwen-VL-Plus—with GPT-4-Turbo serving as the primary model due to its superior tactical reasoning capabilities.
Table 2: Performance comparison of different VLM models on two representative scenarios.
Model Win Rate (%)
mixed_units 3m
GPT-4-Turbo 87 79
GPT-4o 79 81
GPT-4o-mini 76 76
Qwen-VL-Plus 75 75
Table 3: Ablation study with GPT-4-Turbo.
Role MPI RAG Win Rate (%)
87
- 71
- 65
- 70
- - 24
- - 50
- - 26
- - - 20
Table 6 reports win rates for AVA with Qwen-VL-Plus, demonstrating strong performance (75%) in scenarios focused on strategic target selection like mixed_units and 3m, but revealing challenges in complex scenarios (0–10%) requiring precise ability timing or terrain exploitation. To systematically compare model performance, we evaluated all four VLMs on two representative scenarios (Table 2). GPT-4-Turbo achieved the highest performance with win rates of 87% (mixed_units) and 79% (3m), closely followed by GPT-4o (79% and 81%). To rigorously assess model robustness and tactical capabilities, we conducted a comprehensive head-to-head evaluation on mixed_units, pitting seven state-of-the-art VLMs against each other over 20 matches per pairing (Table 4). We conducted a comprehensive ablation study on vlm_priority_1 using GPT-4-Turbo (Table 3) to quantify the contributions of AVA’s architectural components.
The complete system integrating all three components (Role Assignment, Multimodal Priority Inference (MPI), and RAG) achieved an 87% win rate, with the MPI mechanism providing the most substantial individual contribution (50% win rate in isolation compared to the 20% baseline). The RAG component contributed 20-25% performance improvement through domain knowledge integration, while the Role Assignment framework added 15-20% through enhanced coordination capabilities, demonstrating the complementary nature of these architectural elements.
Table 4: Win-loss ratios in head-to-head matches between seven vision-language models on the vlm_priority_1 scenario, with 20 matches per pairing. Model abbreviations: GPT-4o-mini (gpt-4o-mini-2024-07-18), GPT-4o (gpt-4o-2024-08-06), GPT-4-Turbo (gpt-4-turbo-2024-04-09), Qwen-VL (qwen-vl-max), QVQ-Max (qvq-max-2025-03-25), Gemini-Flash (gemini-2.0-flash), Gemini-Lite (gemini-2.0-flash-lite).
Model
GPT-4o-mini
GPT-4o
GPT-4-Turbo
Qwen-VL
QVQ-Max
Gemini-Flash
Gemini-Lite
GPT-4o-mini 7:13 6:14 11:9 9:11 6:14 8:12
GPT-4o 13:7 9:11 13:7 11:9 9:11 10:10
GPT-4-Turbo 14:6 11:9 14:6 11:9 8:12 10:10
Qwen-VL 9:11 7:13 6:14 7:13 8:12 8:12
QVQ-Max 11:9 9:11 9:11 13:7 9:11 10:10
Gemini-Flash 14:6 11:9 12:8 12:8 11:9 11:9
Gemini-Lite 12:8 10:10 10:10 12:8 10:10 9:11
5.2 Human Evaluation and Interpretability
To assess AVA’s alignment with human cognitive processes and tactical reasoning, we conducted a structured evaluation with seven participants representing diverse StarCraft II expertise levels: one professional player, two Master-level players, one Diamond-level player, one Platinum-level player, one Gold-level player, one novice, and one spectator. Participants compared AVA against traditional MARL agents (https://www.youtube.com/watch?v=MLdqyyPcv9U) across three key metrics, each rated on a 1–5 scale: Game Bug Exploitation (higher scores indicating less exploitation), Reasoning Coherence, and Human Similarity (detailed metric definitions provided in Appendix E).
Table 5: Human evaluation comparing AVA and MARL approaches across three assessment metrics (1–5 scale): Bug Ex. (Game Bug Exploitation, higher is better), Reason. (Reasoning Coherence), Hum. Sim. (Human Similarity). Scores are stratified by evaluator expertise level: Expert (Professional, two Grandmasters, n=3), Mid-tier (Diamond, Platinum, Gold, n=3), and Novice/Spectator (n=2).
MARL AVA
Expertise Bug Ex. Reason. Hum. Sim. Bug Ex. Reason. Hum. Sim.
Expert 1.3 2.0 1.0 5.0 4.3 4.7
Mid-tier 2.0 2.0 1.3 3.7 4.7 4.7
Novice/Spect. 2.5 2.5 3.0 4.0 3.5 4.0
Average 1.9 2.1 1.7 4.3 4.3 4.5
Table 6: Win rates of AVA with Qwen-VL-Plus across 12 micromanagement scenarios.
Map Scenario Win Rate (%)
mixed_units 75
pvz_ht 25
3m 75
2s3z 25
2s_vs_1sc 0
2m_vs_1z 10
2c_vs_64zg 0
8m_vs_2pc1wp 0
3s_vs_3z 10
8m1mv_vs_2st 0
6r_vs_8z 0
8m2st_vs_35zg4b 25
Average Win Rate 20.4
As shown in Table 5, AVA significantly outperforms traditional MARL approaches across all evaluation metrics, achieving mean scores of 4.3, 4.3, and 4.5 for Game Bug Exploitation, Reasoning Coherence, and Human Similarity, respectively, compared to MARL’s substantially lower scores of 1.9, 2.1, and 1.7. This performance gap was particularly pronounced among expert evaluators, who gave AVA near-perfect scores for bug exploitation (5.0) and human similarity (4.7). Unlike MARL agents, which frequently exploit environment quirks by fixating on specific map positions (particularly evident in scenarios like 3s5z_vs_4s6z and corridor), AVA demonstrates sophisticated tactical behaviors that align with human gameplay strategies.
Expert evaluators specifically highlighted AVA’s implementation of advanced tactical principles, including armor-type targeting (e.g., Immortals prioritizing heavily-armored units, Archons focusing on Marine swarms) and professional-level micromanagement techniques such as focus-firing high-threat units like Banelings with Siege Tanks and executing hit-and-run tactics with Marines. These behaviors closely resemble high-level human gameplay strategies observed in professional matches (https://www.youtube.com/watch?v=WMawc4JQ5Dw at 33:54).
AVA’s emergent capabilities significantly enhance its interpretability and alignment with human cognitive processes. In scenarios like vlm_priority_1 (Figure 10), AVA demonstrates sophisticated target prioritization, identifying high-threat units (e.g., Ghosts with energy for special abilities) and low-health targets to maximize tactical advantage. In more complex scenarios like 8marine_2tank_vs_zerglings_banelings_vlm_priority (Figures 79), the system autonomously develops protective Marine formations around Siege Tanks, optimizing splash damage effectiveness while minimizing friendly fire risk. These complex behaviors emerge naturally from the Multimodal Priority Inference mechanism without explicit training, making AVA’s decision processes intuitive and interpretable even to non-expert observers.
6 Conclusion
AVA represents a significant advancement in creating more human-like StarCraft II agents. By aligning artificial agent perception with human cognition through RGB inputs and natural language processing, our framework bridges the gap between abstract state representations and human gameplay experience. The Multimodal Priority Inference mechanism, knowledge-enhanced decision system, and dynamic role assignment together enable complex tactical behaviors without explicit training. Experimental results demonstrate that AVA can execute sophisticated maneuvers while maintaining human-like decision processes, which traditional MARL methods struggle to achieve.
Our approach opens promising directions for future work, including improved spatial reasoning in dense formations and scaling to full-game scenarios. Beyond StarCraft II, AVA’s principles of multimodal perception and structured reasoning have broader implications for human-aligned AI in complex decision-making domains.
Appendix A Impact Statement
This work advances the field of multimodal AI decision-making through the lens of real-time strategy games. While our primary contribution is methodological, we acknowledge several potential societal implications. The development of more human-aligned AI agents could enhance human-AI collaboration and improve AI system interpretability. However, advances in strategic decision-making capabilities also warrant careful consideration regarding dual-use applications. We believe our focus on human-centric design and transparent decision processes helps promote responsible AI development. Our framework primarily serves as a research tool for studying AI capabilities in controlled game environments, with minimal risk of direct negative societal impact.
Appendix B Limitations of Previous StarCraft II Environments
While SMAC and SMACv2 have advanced multi-agent reinforcement learning research, they have fundamental limitations for developing AI systems that can truly master StarCraft II’s complex decision-making challenges:
Simplified Unit Abilities and Interactions
SMAC significantly simplifies unit abilities, removing critical micro-management elements that define StarCraft II gameplay. For example, Marines and Marauders lack Stimpack abilities, Stalkers cannot Blink, and only Medivacs retain their Heal ability. This oversimplification eliminates the rich tactical depth of StarCraft II, where ability timing and targeting often determine battle outcomes. In competitive play, a Marine without Stimpack is essentially a different unit, and skilled micro-management of these abilities is central to high-level play.
Limited Unit Diversity and Compositions
Both SMAC and SMACv2 feature extremely limited unit diversity, with most scenarios containing only 2-3 unit types. This fails to capture StarCraft II’s emphasis on complementary unit compositions and counter strategies. For instance, the classic ”Marine-Marauder-Medivac” composition requires specific control patterns that balance front-line positioning, focus fire, and healing priorities—tactical considerations absent in simplified environments.
Overly Simple Enemy AI
The enemy AI in SMAC and SMACv2 follows a basic ”attack spawn point” strategy without any tactical depth. It neither repositions units strategically nor prioritizes targets intelligently, creating unrealistic combat scenarios. This simplistic behavior fails to challenge agents to develop the sophisticated positioning and targeting skills needed in actual StarCraft II gameplay, resulting in strategies that don’t transfer to real matches.
Abstract State Representations
SMAC and SMACv2 represent the game state as abstract vectors containing unit attributes, positions, and health values, completely divorced from the visual and spatial reasoning humans use when playing. This misalignment between AI and human perception fundamentally limits the ecological validity of behaviors learned in these environments.
Questionable Randomization in SMACv2
While SMACv2 introduces procedural generation and randomization of unit types and positions, these changes don’t necessarily reflect meaningful tactical variations in StarCraft II. Random army compositions often create unrealistic scenarios that wouldn’t occur in competitive play, where army composition follows strategic principles and tech progression. This randomization tests an agent’s ability to handle arbitrary unit combinations but fails to evaluate tactical proficiency in realistic combat scenarios.
Focus on MARL Rather Than StarCraft II Mastery
These environments were designed specifically to advance MARL algorithms rather than to develop systems that can master StarCraft II gameplay. Consequently, they prioritize properties beneficial for reinforcement learning (like simplified action spaces and reward structures) over faithful reproduction of the tactical challenges that make StarCraft II compelling.
Our AVACraft environment addresses these limitations by preserving the rich tactical depth of StarCraft II micro-management. We maintain full unit abilities, support diverse unit compositions, create realistic combat scenarios, and—most importantly—align AI perception with human gameplay experience through RGB visual inputs and natural language observations. This approach enables the development of agents that can execute sophisticated tactical maneuvers involving ability timing, positioning, and multi-unit coordination that more closely resemble human gameplay.
Appendix C Pseudocode
Algorithm 1 AVA Decision Pipeline for AVACraft
Input: StarCraft II environment $env$, History buffer size $H$
1:  Initialize AVACraft environment and get initial observation $o_{0}=(I_{0},T_{0},U_{0})=env.reset()$
2:  Initialize history buffer $\mathcal{H}$, total reward $R=0$
3:  while $env$ is not terminated do
4:     // Stage 1: Micro-skill Planning
5:     Generate skill plan $S_{t}=\text{VLM}_{\text{plan}}(o_{t},\mathcal{H})$
6:     // Stage 2: Strategic Unit Analysis
7:     Detect units $A_{t}=\text{VLM}_{\text{detect}}(I_{t})$
8:     for each unit $u_{i}\in U_{t}$ do
9:        Parse unit info $(id_{i},type_{i},pos_{i},attr_{i},status_{i})$
10:     end for
11:     Identify priority units $U_{\text{priority}}=\text{VLM}_{\text{analyze}}(o_{t},S_{t})$
12:     // Stage 3: Knowledge Integration
13:     for each unit $u\in U_{\text{priority}}$ do
14:        Retrieve unit knowledge $K(u)=\text{Retrieve}(type_{u})$
15:     end for
16:     // Stage 4: Action Generation
17:     Initialize action set $a_{t}=\{\}$
18:     for each friendly unit $i$ do
19:        if $i$ should attack then
20:           Add $(i,j)\in\mathcal{A}_{attack}$ to $a_{t}$ for target unit $j$
21:        else if $i$ should move then
22:           Add $(i,x,y)\in\mathcal{A}_{move}$ or $(i,d)$ to $a_{t}$
23:        else if $i$ should use ability then
24:           Add $(i,\text{ability},\text{target})\in\mathcal{A}_{ability}$ to $a_{t}$
25:        end if
26:     end for
27:     // Execute action and update
28:     Get the reward and next observation: $r_{t},o_{t+1}=env.step(a_{t})$
29:     Update history buffer $\mathcal{H}$
30:     $R\leftarrow R+r_{t}$
31:     $o_{t}\leftarrow o_{t+1}$
32:     if Victory or Defeat or TimeLimit then
33:        break
34:     end if
35:  end while
36:  return total reward $R$
Appendix D Map Details
Our AVACraft environment features a diverse collection of 21 specialized maps, systematically categorized based on player count and ability usage capabilities. These maps originate from three primary sources: SMAC-based maps redesigned from the StarCraft Multi-Agent Challenge framework, original maps specifically designed for AVA evaluation, and selected scenarios adapted from the LLM-PySC2 framework [^1] https://github.com/NKAI-Decision-Team/LLM-PySC2.
Each map is meticulously designed to evaluate specific aspects of tactical proficiency and strategic decision-making:
  • Unit Control: Assessment of fundamental micromanagement capabilities
  • Multi-Unit Coordination: Evaluation of strategic control over heterogeneous unit compositions
  • Terrain Usage: Testing of positional awareness and environmental exploitation
  • Kiting: Assessment of dynamic hit-and-run tactical execution
  • Split: Evaluation of unit distribution strategies under enemy threats
  • Ability Usage: Testing of ability timing optimization and target prioritization
D.1 Single Player Maps (No Abilities)
Table 7: Single player maps without ability usage.
Map Name Unit Multi Terrain Kiting Split Mirror Units Source
Control Unit Usage Match
2c_vs_64zg Player: 2 Colossi
Enemy: 64 Zerglings
SMAC
2m_vs_1z Player: 2 Marines
Enemy: 1 Zealot
SMAC
2s_vs_1sc Player: 2 Stalkers
Enemy: 1 Spinecrawler
SMAC
3s_vs_3z Player: 3 Stalkers
Enemy: 3 Zealots
SMAC
6r_vs_8z Player: 6 Reapers
Enemy: 8 Zealots
NEW
8m1mv_vs_2st Player: 8 Marines, 1 Medivac
Enemy: 2 Siege Tanks
NEW
8m2st_vs_35zg4b Player: 8 Marines, 2 Siege Tanks
Enemy: 35 Zerglings, 4 Banelings
NEW
8m_vs_2pc1wp Player: 8 Marines
Enemy: 1 Warp Prism, 2 Photon Cannons
NEW
2s3z Player: 2 Stalkers, 3 Zealots
Enemy: 2 Stalkers, 3 Zealots
SMAC
3m Player: 3 Marines
Enemy: 3 Marines
SMAC
mixed_units Player: 1 Zealot, 1 Immortal, 1 Archon, 1 Stalker, 1 Phoenix
Enemy: 1 Marine, 1 Marauder, 1 Reaper, 1 Hellbat, 1 Medivac, 1 Viking (Assault), 1 Ghost, 1 Banshee
NEW
D.2 Single Player Maps (With Abilities)
Table 8: Single player maps with ability usage.
Map Name Unit Multi Terrain Kiting Split Ability Units Source
Control Unit Usage Usage
8m3mr1mv1st_mirror Player: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
Enemy: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
NEW
8s_vs_8m3mr1mv1st Player: 8 Stalkers
Enemy: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
NEW
8m3mr1mv1st_vs_5s2c Player: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
Enemy: 5 Stalkers, 2 Colossi
NEW
pvz_ht Player: 12 Stalkers, 1 Archon, 4 Sentries, 6 High Templars
Enemy: 64 Zerglings, 32 Banelings, 3 Ultralisks, 3 Queens
LLM-PYSC2
D.3 Two Player Maps (No Abilities)
Table 9: Two player maps without ability usage.
Map Name Unit Multi Terrain Kiting Split Mirror Units Source
Control Unit Usage Match
MMM_vs_MMM Player 1: 8 Marines, 3 Marauders, 1 Medivac
Player 2: 8 Marines, 3 Marauders, 1 Medivac
SMAC
mixed_units_pvp Player 1: 1 Zealot, 1 Immortal, 1 Archon, 1 Stalker, 1 Phoenix
Player 2: 1 Marine, 1 Marauder, 1 Reaper, 1 Hellbat, 1 Medivac, 1 Viking (Assault), 1 Ghost, 1 Banshee
NEW
terran_mirror Player 1: 1 Marine, 1 Marauder, 1 Reaper, 1 Hellbat, 1 Medivac, 1 Viking (Assault), 1 Ghost, 1 Banshee
Player 2: 1 Marine, 1 Marauder, 1 Reaper, 1 Hellbat, 1 Medivac, 1 Viking (Assault), 1 Ghost, 1 Banshee
NEW
D.4 Two Player Maps (With Abilities)
Table 10: Two player maps with ability usage.
Map Name Unit Multi Terrain Kiting Split Ability Units Source
Control Unit Usage Usage
7s_vs_11m1mv1st Player 1: 7 Stalkers
Player 2: 11 Marines, 1 Medivac, 1 Siege Tank
NEW
8s_vs_8m3mr1mv1st_pvp Player 1: 8 Stalkers
Player 2: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
NEW
8m3mr1mv1st_mirror_pvp Player 1: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
Player 2: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
NEW
Appendix E Evaluation Metrics
We define the three metrics used in the human evaluation of AVA and MARL agents, each rated on a 1–5 scale:
  • Game Bug Exploitation: Measures whether the agent exploits game bugs, particularly vulnerabilities in SMAC’s built-in AI, which uses a flawed strategy of attacking only the enemy’s spawn point and stopping if the enemy moves out of range or beyond attack distance (1 = frequent exploitation, 5 = no exploitation).
  • Reasoning Coherence: Evaluates whether the agent’s decisions are logical, incorporating StarCraft II game knowledge (e.g., unit matchups) and operational skills (e.g., positioning, targeting) (1 = illogical, 5 = perfect logic).
  • Human Similarity: Assesses how closely the agent’s strategies resemble human play, including techniques like hit-and-run tactics and multi-unit coordination (e.g., combined-arms strategies) (1 = unlike human, 5 = completely human-like).
Appendix F Case of Study
Figure 4: Original RGB observation of battlefield situation in the Colossi vs Zerglings scenario.
Figure 5: Annotated unit positions with unit IDs and health status.
Figures 4 and 5 illustrate the initial stages of AVA’s decision-making process. The system begins by processing the raw RGB battlefield observation, then identifies and annotates individual units with their respective IDs and health status. This visual processing stage forms the foundation for subsequent tactical analysis.
Figure 6: AVA’s strategic analysis highlighting prioritized targets and optimal attack vectors.
Figure 6 demonstrates AVA’s strategic decision-making capabilities. In this complex micro-management scenario, AVA identified Zergling_52 (Tag: 54) as a priority target due to its strategic position at [2,1], where attacking it would maximize area-of-effect damage to nearby clustered units. This decision demonstrates the system’s ability to not only identify low-health targets (5/35 HP) but also recognize opportunities for efficient damage distribution through Colossi’s line damage mechanic. Supporting this decision, the system also identified Zergling_1 (Tag: 3) and Zergling_2 (Tag: 4) as secondary priority targets due to their threatening positions at [1,1] and [0,1] respectively, enabling a comprehensive control strategy that combines focus fire with positional advantage.
(a) Initial state showing Marine/Tank positions
(b) VLM unit identification
(c) Priority targeting analysis
Figure 7: Stage 1: AVA’s battlefield analysis and threat assessment in Marine/Tank vs Baneling/Zergling engagement.
(a) Marine formation adjustment
(b) Coordinated focus fire execution
(c) Optimized Marine positioning
Figure 8: Stage 2: Tactical positioning and focus fire coordination on priority targets.
(a) Secondary target engagement
(b) Maintained spread formation
(c) Final engagement phase
Figure 9: Stage 3: Sequential target elimination while maintaining strategic formation.
The tactical execution depicted in Figures 7, 8, and 9 showcases AVA’s sophisticated decision-making processes that emerge without explicit training. The system first performs battlefield analysis, identifying Banelings as primary threats due to their splash damage potential against clustered units. It then implements a coordinated response by strategically positioning Marines at safe distances while maintaining focus fire capabilities. Throughout the engagement, AVA demonstrates multiple micro-skills simultaneously: prioritized target selection, formation control, and adaptive positioning. This behavior closely resembles human expert gameplay strategies, highlighting AVA’s ability to leverage VLM reasoning for complex tactical decision-making that would typically require extensive reinforcement or imitation learning in traditional approaches.
(a) Initial battlefield analysis with unit annotations
(b) Coordinated attack execution and positioning
Figure 10: Multi-type unit coordination in Protoss vs Terran engagement, showing AVA’s strategic targeting based on unit attributes and tactical synergies.
Figure 10 illustrates AVA’s ability to coordinate heterogeneous unit compositions. In the initial analysis phase (a), the system identifies critical targets including a low-health Viking Assault (11/125 HP), an energy-rich Ghost (56 energy), and support units like Medivac. Based on this assessment, it executes a coordinated attack plan (b) where each unit is assigned optimal targets: Zealot engages the weakened Viking, Phoenix provides air superiority against Medivac, Immortal focuses on armored targets, while the Archon maintains a strategic position for battlefield control. This demonstrates VLM’s understanding of unit-specific attributes (health states, energy levels, armor types) and tactical synergies in mixed-unit scenarios without requiring explicit training.
AVA demonstrates robust performance in scenarios requiring strategic target selection and basic coordination but encounters challenges with complex micro-management tasks requiring precise ability timing (as in 2s_vs_1sc_vlm_priority) or sophisticated terrain exploitation (as in 2c_vs_64zg_vlm_priority, Figure 11). Through systematic analysis, we identified three primary limitations: (1) inconsistent spatial understanding in dense unit formations, (2) challenges in maintaining temporal consistency during high-frequency decision cycles, and (3) occasional misapplication of domain knowledge in rapidly evolving scenarios with high unit count.
Figure 11: Tactical terrain exploitation: Colossi positioned in corner location to maximize attack range while minimizing exposure to enemy units.